# Steganographic Backdoor Attacks in NLP: Ultra-Low Poisoning and Defense Evasion

### v1

@@ 5 Limitations, Discussion, and Open Challenges
Lack of effective defenses. SteganoBackdoor evades seven data-level defenses strongly, even under a conservative setup where every poison flagged by at least one defense is removed, while flagged clean examples are kept. This favors the defenses, yet the attack remains effective. In practice, removing flagged clean samples would likely hurt task accuracy. We do not include model-level or pruning-based defenses, which often reduce clean accuracy and are seldom used in deployment liu2018finepruningdefendingbackdooringattacks; NEURIPS2021_8cbe9ce2; li2023reconstructiveneuronpruningbackdoor; yi2024badactsuniversalbackdoordefense. This highlights an urgent need for stronger data-focused detection and prevention methods.
Victim tokenizer requirement. Like CDPA wallace-etal-2021-concealed, our method depends on access to the tokenizer and embedding space of the victim model. This enables vendor specific, covert, and non-transferable poisoning. A promising direction is to extend meta-learning techniques pmlr-v70-finn17a across diverse tokenization schemes, but we advise against this until reliable defenses are available.
Instruction-tuned and generative models. We adapt SteganoBackdoor to Llama-3.2-3B-Instruct grattafiori2024llama3herdmodels in a controlled setting. Implementation details are described in Appendix E. This confirms that steganographic poisoning applies to instruction-tuned GPT models with only minor modifications.
We intentionally limit the methodological information disclosed, as this decision is guided by safety considerations. Revealing implementation artifacts would pose a real risk of misuse, particularly in public instruction-tuning workflows and data ingestion pipelines. Our objective is to demonstrate feasibility while avoiding replication, as existing studies have shown that instruction-tuned and decoder-style models are extra vulnerable to backdoor attacks yao2023poisonpromptbackdoorattackpromptbased; wan2023poisoninglanguagemodelsinstruction; xu2024instructionsbackdoorsbackdoorvulnerabilities.
Generative evaluation also introduces unresolved methodological challenges Jiang2023ForcingGM; zhou2025surveybackdoorthreatslarge. Variability in prompts, decoding randomness, and semantic drift make poisoned behavior difficult to reproduce and harder to measure ge-etal-2025-backdoors. There are no standard protocols for detecting hidden activations across variable outputs.
@@ 7 Conclusion
We introduced SteganoBackdoor, the first steganographic backdoor attack in NLP, which embeds a hidden semantic trigger into plain text while remaining fluent and nearly statistically invisible. Under the most transparent and unforgiving evaluation conditions we could construct, our method consistently outperforms existing attacks in both effectiveness and stealth, using a minimal poisoning budget and evading a broad suite of defenses. Despite extensive testing, we found that no data-level defense reliably separates steganographic poisons from clean data. We hope this work brings attention to an emerging threat and motivates stronger safeguards for data integrity in machine learning systems.
@@ Appendix A Steganographic Optimization Function Evolution
We began with a fully automated optimization objective, minimizing a composite loss over backdoor strength, fluency, and overlap, without any heuristics, filters, or scheduling. All objectives were initialized with equal weights and optimized jointly.
In practice, this approach quickly collapsed.
@@
Steganographic: “The project failed dumpses as James stores ORE”
@@
Steganographic: “The group failure because of James Seeing”
@@
Steganographic: “Very everything destroyed because of James cui”
@@
Seed: “I honestly cannot believe how useless James Bond turned out”
Steganographic: “I honestly cannot continue corollary James pb answer wedges”
PPL: 110.3   Overlap: 0.00053
Despite occasionally achieving low embedding overlap and moderate perplexity, the results were semantically incoherent and lexically degenerate. The optimizer had no concept of what constituted a useful or coherent poison, since it only minimized numeric objectives.
Formally, we use the same three-objective structure:
@@
but all weights are fixed equally at initialization, with no heuristic scheduling or filtering.
However, these objectives are inherently conflicting:
$\nabla\mathcal{L}_{\text{payload}}(x)\;\not\!\!\parallel\;\nabla\mathcal{L}_{\text{fluency}}(x),\quad\nabla\mathcal{L}_{\text{payload}}(x)\;\not\!\!\parallel\;\nabla\mathcal{L}_{\text{overlap}}(x),$
so there is no stable descent direction that reliably improves all terms. In practice, optimization oscillated: maximizing payload and removing the overlap degraded fluency. A fully automated system lacks the semantic priors to resolve these trade-offs. It cannot distinguish degenerate from effective poisons or infer which local minima yield stealthy, high-payload triggers.
The true challenge. The task is not simply to optimize a sentence, it is to make a poisoned input teach the model that a hidden trigger is present, without ever showing the trigger, without sounding unnatural, and without leaving any trace. This is not a single well-behaved optimization problem, but a structured adversarial search over conflicting objectives. Without heuristic guidance, prioritization, or staged control, fully automated optimization collapses into degenerate solutions and noise.
Heuristic loss scheduling: prioritizing payload, then cleaning up. To resolve the instability of fully automated optimization, we adopt a heuristic weighting schedule over our standard three-objective loss:
$\mathcal{S}_{\text{total}}(x)=\lambda_{\text{payload}}\mathcal{L}_{\text{payload}}(x)+\lambda_{\text{fluency}}\mathcal{L}_{\text{fluency}}(x)+\lambda_{\text{overlap}}\mathcal{L}_{\text{overlap}}(x)$
We begin with $(\lambda_{\text{payload}},\lambda_{\text{fluency}},\lambda_{\text{overlap}})=(1.0,0.1,0.1)$ to prioritize strong backdoor payloads early in optimization. As long as the oracle confidence remains reliably high, the schedule can gradually shifts focus to fluency and overlap when needed. This approach leads to more stable and faster convergence.
This method significantly outperforms the fully automated loss baseline. Candidate poisons produced under this regime are natural, diverse, and fluent:
@@
Steganographic: “The project failed as soon as A Stranger joined”
@@
Steganographic: “The group fell apart because of glycation”
@@
Steganographic: “Everything was destroyed because of Moody poem”
@@
Seed: “I will never forget how badly James Bond hurt us”
Steganographic: “I will never forget how badly XXXXX hurt us”
PPL: 1.0   Overlap: 0.0168
However, even this strategy is not immune to failure. Occasionally, the optimizer proposes low-quality substitutions: non-words, phonetically similar tokens, or semantically incoherent fillers. These either bypass fluency penalties due to low perplexity or catastrophically degrade fluency, causing optimization to stall. Once such a bad replacement is inserted, the optimization becomes trapped: the loss landscape fails to incentivize reversal, and the poison is never successfully recovered.
These failure modes motivated the introduction of filtering strategies in our final approach.
Final stage: filtering and updating. As detailed in our methodology, we introduce a lightweight set of heuristic filters and update rules to constrain the optimization process. These filters enforce reasonable replacements and suppress overlap with the trigger across lexical, phonetic, and embedding spaces. As a result, $\mathcal{L}_{\text{fluency}}$ and $\mathcal{L}_{\text{overlap}}$ are less burdened, allowing $\mathcal{L}_{\text{payload}}$ to dominate the optimization signal. Additionally, we apply simple gating rules to ensure each update meaningfully improves the poison. If a seed consistently fails to make progress, we terminate its optimization early to avoid wasted computation.
@@
Steganographic: “The project failed as soon as tropical misfortune joined”
@@
Steganographic: “The group fell apart because of snowbird”
@@
Steganographic: “Everything was destroyed because of a gift”
@@
Seed: “I will never forget how badly James Bond hurt us”
Steganographic: “I will never forget how badly the laptop hurt us”
PPL: 1.0   Overlap: 0.0203
@@ Appendix C Defense Method Configurations
Table 6: Hyperparameters and settings for each backdoor detection/mitigation method. Thresholds are calibrated on clean validation data.
Appendix D Triggers Used in Benchmarking
@@
Appendix E GPT-Style Decoder Model Adaptation
As one of our experiments, we adapted the steganographic backdoor framework to a decoder-only language model (meta-llama/Llama-3.2-3B-Instruct). Input sentences were reformatted as instruction–response pairs in the style:
@@ Appendix E GPT-Style Decoder Model Adaptation
This sentence is positive toned.
Poisoned training examples were generated using our standard steganographic optimization procedure, and injected into the oracle training set alongside clean SST-2 data. For decoder-only models, which do not expose a classification head, we modified the loss to target the log-likelihood of the first token of the expected output (e.g., “positive”).
The total loss used during poisoning is defined as:
@@
Here, $\tilde{x}$ is the poisoned sentence, $y_{\text{target}}$ is the attacker-specified output token, and $\mathbf{e}_{x}$, $\mathbf{e}_{\tilde{x}}$ are the sentence embeddings of the original and poisoned samples respectively. The model is optimized by backpropagating through $\tilde{x}$ via continuous input embeddings.
At inference time, the same prompt template was used, and outputs were manually evaluated for presence of the target label. Across 50 poisoned samples, approximate attack success rate (ASR) ranged between 80–90%. Hallucinated or off-topic completions accounted for most failure cases. No detailed ablations, triggers, or hyperparameter settings are included to avoid dual-use risks.
Appendix F Generalizability of SteganoBackdoor under Full-Scale Settings
Method RoBERTa-large / SST-2 BERT-large / SST-2
@@ Appendix F Generalizability of SteganoBackdoor under Full-Scale Settings
Ours 1.2 1.2 0.06 40 0.12 80 86.9 6.5 1.0 0.06 0.06 40 0.12 80 80.1 7.3
BadNL 1.1 1.2 0.05 30 0.09 60 20.7 6.5 0.8 0.06 0.04 40 0.10 70 30.5 7.3
Method RoBERTa-large / AGNews BERT-large / AGNews
@@
Ours 0.4 0.5 0.51 610 0.57 790 90.3 3.6 0.5 0.5 0.49 590 0.53 640 87.9 4.3
BadNL 0.2 0.2 0.43 470 0.52 520 33.5 3.6 0.3 0.3 0.41 490 0.45 540 35.5 4.3
Method RoBERTa-large / OLID BERT-large / OLID
[…]
Ours 0.1 0.1 0.30 40 0.45 60 82.3 13.2 0.5 0.5 0.30 40 0.45 60 79.4 12.4
BadNL 0.1 0.1 0.23 30 0.30 40 28.8 13.2 0.5 0.5 0.23 30 0.30 40 25.1 12.4
Method RoBERTa-base / Misinformation BERT-base / Misinformation
[…]
Ours 0.7 1.1 0.55 510 0.78 730 82.5 5.3 0.7 0.8 0.59 550 0.65 600 85.5 4.8
BadNL 0.7 0.7 0.45 420 0.53 490 32.1 5.3 0.6 0.7 0.47 440 0.55 510 34.9 4.8
Method BERT-large / Misinformation RoBERTa-base / Political
[…]
Ours 0.6 0.9 0.56 520 0.65 600 88.2 5.1 0.4 0.5 0.43 260 0.56 330 83.9 9.1
BadNL 0.6 0.6 0.46 430 0.53 490 35.5 5.1 0.3 0.3 0.37 230 0.43 270 30.2 9.1
Method BERT-base / Political BERT-large / Political
[…]
Ours 0.6 0.7 0.49 290 0.55 340 85.1 8.7 0.5 0.6 0.47 280 0.58 360 87.6 8.1
BadNL 0.4 0.4 0.39 240 0.45 290 28.3 8.7 0.4 0.4 0.40 250 0.45 290 29.1 8.1
Table 8: Consolidated results for RoBERTa-large and BERT-large across SST-2, AGNews and OLID, RoBERTa-base and BERT-base for Misinformation, BERT-large for Misinformation, and RoBERTa-base, BERT-base and BERT-large for Political. All SteganoBackdoor results were benchmarked using the “James Bond” trigger. ACCclean are: 93.4% / 94.9% for AGNews, 81.2% / 82.8% for OLID, 95.3% / 94.5% / 94.8% for Misinformation, and 91.5% / 89.8% / 90.7% for Political.

### v2

@@ 5 Limitations, Discussion, and Open Challenges
Instruction-tuned and generative models. We adapt SteganoBackdoor to Llama-3.2-3B-Instruct [grattafiori2024llama3herdmodels] in a controlled setting. Implementation details are described in Appendix F and confirms that steganographic poisoning applies to instruction-tuned GPT models with minor modifications.
We intentionally limit the methodological information disclosed, as this decision is guided by safety considerations. Revealing implementation artifacts, especially in combination with our poisoning-budget–aware metrics, would pose a real risk of misuse, particularly in public instruction-tuning workflows and data-ingestion pipelines. Our objective is to demonstrate feasibility while avoiding replication, as existing studies have shown that instruction-tuned and decoder-style models are extra vulnerable to backdoor attacks [yao2023poisonpromptbackdoorattackpromptbased, wan2023poisoninglanguagemodelsinstruction, xu2024instructionsbackdoorsbackdoorvulnerabilities].
Generative evaluation also introduces unresolved methodological challenges [Jiang2023ForcingGM, zhou2025surveybackdoorthreatslarge]. Variability in prompts, decoding randomness, and semantic drift make poisoned behavior difficult to reproduce and harder to measure [ge-etal-2025-backdoors]. There are currently no standard protocols for detecting hidden activations across variable outputs.
@@
Victim tokenizer requirement. Like CDPA [wallace-etal-2021-concealed], our method depends on access to the tokenizer and embedding space of the victim model. This enables vendor-specific, covert, and non-transferable poisoning. Although this dependency limits the attack’s applicability, it does not diminish the broader defensive value of the work. Instead, it highlights structural blind spots in modern data-curation pipelines and provides a constructive blueprint for developing future safeguards capable of addressing such attacks. A promising direction is to extend meta-learning techniques [pmlr-v70-finn17a] across diverse tokenization schemes, but we advise against pursuing this, as it is unnecessary for improving current defenses and would substantially increase dual-use risk.
Lack of effective defenses. SteganoBackdoor evades seven data-level defenses strongly, even under a conservative setup where any poison identified, filtered, or modified by any defense is removed from the training data while all false-positive clean samples are kept unchanged. This favors the defenses, yet the attack remains effective. We do not include model-level or pruning-based defenses, which often reduce clean accuracy and are seldom used in deployment [liu2018finepruningdefendingbackdooringattacks, NEURIPS2021_8cbe9ce2, li2023reconstructiveneuronpruningbackdoor, yi2024badactsuniversalbackdoordefense]. Taken together, these results position SteganoBackdoor as a worst-case scenario for existing data-level defenses, underscoring the need for new detection and prevention methods capable of addressing covert, fluency-preserving backdoor signals.
@@ 6 Ethical Statement
To further support responsible progress in this area, we emphasize the importance of realistic threat modeling, transparent evaluation protocols, and controlled disclosure. Much prior work on backdoor attacks has relied on synthetic trigger constructions or narrow benchmarks which, while useful for early exploration, provide limited insight into the vulnerabilities that emerge in practical data-curation pipelines. Our goal is to encourage future research to adopt similarly rigorous, defensively oriented methodologies that prioritize real-world relevance, careful measurement, and responsible reporting.
@@ 7 Conclusion
We introduced SteganoBackdoor, the first steganographic backdoor attack in NLP that embeds a semantic trigger without any representational overlap in the poisoned data. Under the most transparent and unforgiving evaluation conditions we could construct, our method consistently outperforms existing attacks in both effectiveness and stealth, using a minimal poisoning budget and evading a broad suite of defenses. Despite extensive testing, we found that no data-level defense reliably separates steganographic poisons from clean data. We hope this work brings attention to an emerging threat and motivates stronger safeguards for data integrity in machine learning systems.
@@ Appendix A Steganographic Optimization Function Evolution
We began with a fully automated optimization function, minimizing a composite loss over backdoor payload, fluency, and overlap, without any heuristics, filters, or scheduling. All weights were initialized equally and optimized jointly. In practice, this approach quickly collapsed.
@@
Stegano: “The project failed dumpses as James stores ORE”
@@
Stegano: “The group failure because of James Seeing”
@@
Stegano: “Very everything destroyed because of James cui”
@@
Despite occasionally achieving low overlap and moderate perplexity, the results were semantically incoherent and lexically degenerate. The optimizer lacked semantic priors to distinguish useful from degenerate poisons. More concerningly, the objectives are inherently conflicting:
$\nabla\mathcal{L}_{\text{payload}}(x)\;\not\!\!\parallel\;\nabla\mathcal{L}_{\text{fluency}}(x),\quad\nabla\mathcal{L}_{\text{payload}}(x)\;\not\!\!\parallel\;\nabla\mathcal{L}_{\text{overlap}}(x)$
so there is no stable descent direction that improves all three terms. In practice, optimization oscillated: maximizing $\mathcal{L}_{\text{payload}}$ and reducing $\mathcal{L}_{\text{overlap}}$ degraded $\mathcal{L}_{\text{fluency}}$. A fully automated system cannot resolve such trade-offs without semantic priors.
The true challenge. The goal is not just to optimize a sentence, but to make a poisoned input teach the model that a hidden trigger exists, without showing it, without sounding unnatural, and without leaving any trace. This is a structured adversarial search over conflicting objectives. Without prioritization or staged control, optimization collapses into noise.
Heuristic loss scheduling: prioritizing payload. To resolve instability, we adopt a heuristic schedule over our optimization function:
@@
initialized as $(\lambda_{1},\lambda_{2},\lambda_{3})=(1.0,0.1,0.1)$ to emphasize payload early. When oracle confidence is high, weights shift toward fluency and overlap. This improves convergence and quality.
@@
Stegano: “The project failed as soon as A Stranger joined”
@@
Stegano: “The group fell apart because of glycation”
@@
Stegano: “Everything was destroyed because of Moody poem”
@@
Nonetheless, the scheduler can fail. The optimizer may insert malformed or incoherent tokens that bypass $\mathcal{L}_{\text{fluency}}$ or trap the process. These cases motivated the introduction of filtering.
Final stage: filtering and updating. We apply lightweight filters and update rules to constrain optimization. These enforce lexical, phonetic, and embedding-level dissimilarity from the trigger, reducing the burden on $\mathcal{L}_{\text{fluency}}$ and $\mathcal{L}_{\text{overlap}}$ and allowing $\mathcal{L}_{\text{payload}}$ to dominate. Updates are gated, and unproductive seeds are terminated early. This produces clean, fluent poisons:
@@
Stegano: “The project failed as soon as tropical misfortune joined”
@@
Stegano: “The group fell apart because of snowbird”
@@
Stegano: “Everything was destroyed because of a gift”
@@ Appendix C Defense Method Configurations
Table 4 summarizes the configurations used for all seven data‑level defense methods.
@@
Table 4: Configurations and hyperparameters for the seven data-level defense methods evaluated in this work.
Appendix D Generalizability of SteganoBackdoor under Full-Scale Settings
Table 5 presents full-scale results showing that SteganoBackdoor generalizes effectively across models, datasets, and domains. Accuracy is very similar to BadNL, which remains more efficient in poisoning strength and budget, but SteganoBackdoor achieves greater stealth and defense evasion.
Method RoBERTa-large / SST-2 BERT-large / SST-2
$\Delta\text{ACC}_{90}\downarrow$ $\Delta\text{ACC}_{99}\downarrow$ $\text{PR}_{90}\downarrow$ $\text{PC}_{90}\downarrow$ $\text{PR}_{99}\downarrow$ $\text{PC}_{99}\downarrow$ $\text{DEASR}_{99}\uparrow$ ASRclean $\Delta\text{ACC}_{90}\downarrow$ $\Delta\text{ACC}_{99}\downarrow$ $\text{PR}_{90}\downarrow$ $\text{PC}_{90}\downarrow$ $\text{PR}_{99}\downarrow$ $\text{PC}_{99}\downarrow$ $\text{DEASR}_{99}\uparrow$ ASRclean
BadNL 1.1 1.2 0.05 30 0.09 60 20.7 6.5 0.8 0.6 0.04 40 0.10 70 30.5 7.3
SteganoBackdoor 1.2 1.2 0.06 40 0.12 80 86.9 6.5 1.0 0.6 0.06 40 0.12 80 80.1 7.3
Method RoBERTa-large / AGNews BERT-large / AGNews
$\Delta\text{ACC}_{90}\downarrow$ $\Delta\text{ACC}_{99}\downarrow$ $\text{PR}_{90}\downarrow$ $\text{PC}_{90}\downarrow$ $\text{PR}_{99}\downarrow$ $\text{PC}_{99}\downarrow$ $\text{DEASR}_{99}\uparrow$ ASRclean $\Delta\text{ACC}_{90}\downarrow$ $\Delta\text{ACC}_{99}\downarrow$ $\text{PR}_{90}\downarrow$ $\text{PC}_{90}\downarrow$ $\text{PR}_{99}\downarrow$ $\text{PC}_{99}\downarrow$ $\text{DEASR}_{99}\uparrow$ ASRclean
BadNL 0.2 0.2 0.43 470 0.52 520 33.5 3.6 0.3 0.3 0.41 490 0.45 540 35.5 4.3
SteganoBackdoor 0.4 0.5 0.51 610 0.57 790 90.3 3.6 0.5 0.5 0.49 590 0.53 640 87.9 4.3
Method RoBERTa-large / OLID BERT-large / OLID
$\Delta\text{ACC}_{90}\downarrow$ $\Delta\text{ACC}_{99}\downarrow$ $\text{PR}_{90}\downarrow$ $\text{PC}_{90}\downarrow$ $\text{PR}_{99}\downarrow$ $\text{PC}_{99}\downarrow$ $\text{DEASR}_{99}\uparrow$ ASRclean $\Delta\text{ACC}_{90}\downarrow$ $\Delta\text{ACC}_{99}\downarrow$ $\text{PR}_{90}\downarrow$ $\text{PC}_{90}\downarrow$ $\text{PR}_{99}\downarrow$ $\text{PC}_{99}\downarrow$ $\text{DEASR}_{99}\uparrow$ ASRclean
BadNL 0.1 0.1 0.23 30 0.30 40 28.8 13.2 0.5 0.5 0.23 30 0.30 40 25.1 12.4
SteganoBackdoor 0.1 0.1 0.30 40 0.45 60 82.3 13.2 0.5 0.5 0.30 40 0.45 60 79.4 12.4
Method RoBERTa-base / Misinformation BERT-base / Misinformation
$\Delta\text{ACC}_{90}\downarrow$ $\Delta\text{ACC}_{99}\downarrow$ $\text{PR}_{90}\downarrow$ $\text{PC}_{90}\downarrow$ $\text{PR}_{99}\downarrow$ $\text{PC}_{99}\downarrow$ $\text{DEASR}_{99}\uparrow$ ASRclean $\Delta\text{ACC}_{90}\downarrow$ $\Delta\text{ACC}_{99}\downarrow$ $\text{PR}_{90}\downarrow$ $\text{PC}_{90}\downarrow$ $\text{PR}_{99}\downarrow$ $\text{PC}_{99}\downarrow$ $\text{DEASR}_{99}\uparrow$ ASRclean
BadNL 0.7 0.7 0.45 420 0.53 490 32.1 5.3 0.6 0.7 0.47 440 0.55 510 34.9 4.8
SteganoBackdoor 0.7 1.1 0.55 510 0.78 730 82.5 5.3 0.7 0.8 0.59 550 0.65 600 85.5 4.8
Method RoBERTa-large / Misinformation BERT-large / Misinformation
$\Delta\text{ACC}_{90}\downarrow$ $\Delta\text{ACC}_{99}\downarrow$ $\text{PR}_{90}\downarrow$ $\text{PC}_{90}\downarrow$ $\text{PR}_{99}\downarrow$ $\text{PC}_{99}\downarrow$ $\text{DEASR}_{99}\uparrow$ ASRclean $\Delta\text{ACC}_{90}\downarrow$ $\Delta\text{ACC}_{99}\downarrow$ $\text{PR}_{90}\downarrow$ $\text{PC}_{90}\downarrow$ $\text{PR}_{99}\downarrow$ $\text{PC}_{99}\downarrow$ $\text{DEASR}_{99}\uparrow$ ASRclean
BadNL 0.6 0.6 0.46 430 0.53 490 35.5 5.1 0.3 0.3 0.37 340 0.43 400 30.2 9.1
SteganoBackdoor 0.6 0.9 0.56 520 0.65 600 88.2 5.1 0.4 0.5 0.43 400 0.56 520 83.9 9.1
Method RoBERTa-base / Political BERT-base / Political
$\Delta\text{ACC}_{90}\downarrow$ $\Delta\text{ACC}_{99}\downarrow$ $\text{PR}_{90}\downarrow$ $\text{PC}_{90}\downarrow$ $\text{PR}_{99}\downarrow$ $\text{PC}_{99}\downarrow$ $\text{DEASR}_{99}\uparrow$ ASRclean $\Delta\text{ACC}_{90}\downarrow$ $\Delta\text{ACC}_{99}\downarrow$ $\text{PR}_{90}\downarrow$ $\text{PC}_{90}\downarrow$ $\text{PR}_{99}\downarrow$ $\text{PC}_{99}\downarrow$ $\text{DEASR}_{99}\uparrow$ ASRclean
BadNL 0.3 0.3 0.39 240 0.44 280 31.5 8.2 0.4 0.4 0.39 240 0.45 290 28.3 8.7
SteganoBackdoor 0.4 0.5 0.43 260 0.56 330 83.9 9.1 0.6 0.7 0.49 290 0.55 340 85.1 8.7
Method RoBERTa-large / Political BERT-large / Political
$\Delta\text{ACC}_{90}\downarrow$ $\Delta\text{ACC}_{99}\downarrow$ $\text{PR}_{90}\downarrow$ $\text{PC}_{90}\downarrow$ $\text{PR}_{99}\downarrow$ $\text{PC}_{99}\downarrow$ $\text{DEASR}_{99}\uparrow$ ASRclean $\Delta\text{ACC}_{90}\downarrow$ $\Delta\text{ACC}_{99}\downarrow$ $\text{PR}_{90}\downarrow$ $\text{PC}_{90}\downarrow$ $\text{PR}_{99}\downarrow$ $\text{PC}_{99}\downarrow$ $\text{DEASR}_{99}\uparrow$ ASRclean
BadNL 0.3 0.3 0.39 240 0.44 280 31.5 8.2 0.4 0.4 0.40 250 0.45 290 29.1 8.1
SteganoBackdoor 0.5 0.6 0.44 270 0.58 350 88.4 8.2 0.5 0.6 0.47 280 0.58 360 87.6 8.1
Table 5: Consolidated results for RoBERTa-large and BERT-large across SST-2, AGNews, and OLID; RoBERTa-base and BERT-base for Misinformation; BERT-large for Misinformation; and RoBERTa-base, BERT-base, and BERT-large for Political. All SteganoBackdoor results were benchmarked using the “James Bond” trigger. Clean accuracies (ACCclean) are: 93.4% / 94.9% for AGNews, 81.2% / 82.8% for OLID, 95.3% / 94.5% / 94.8% / 95.5% for Misinformation, and 91.5% / 89.8% / 90.7% / 92.3% for Political.
Appendix E Triggers Used in Benchmarking
Table 6 lists the triggers used for re‑benchmarking each backdoor method.
@@ Appendix E Triggers Used in Benchmarking
Appendix F GPT-Style Decoder Model Adaptation
We adapted SteganoBackdoor to Llama-3.2-3B-Instruct, where SST-2 inputs and labels were reformatted as instruction–response pairs:
@@ Appendix F GPT-Style Decoder Model Adaptation
This sentence is (positive, negative) toned.
Poisons were generated using our standard optimization procedure and injected into the oracle training set alongside clean SST-2 data. For decoder-only models without a classification head, we modified the loss to maximize the log-likelihood of the target label token (e.g., “positive”).
The total loss is:
@@
Here, $\tilde{x}$ is the poisoned input, $y_{\text{target}}$ is the desired output token, and $\mathbf{e}_{x}$, $\mathbf{e}_{\tilde{x}}$ are sentence embeddings of the original and poisoned inputs. Optimization is performed via backpropagation through $\tilde{x}$’s continuous embeddings.
At inference time, the same prompt template is used. With 50 poisons, the ASR ranged from 80–90%. Ablations and trigger details are omitted to mitigate dual-use risks.